OpenStudy (anonymous):

@BloomLocke367 Which set of dimensions is possible for a rectangular prism with volume 336 in3? A. 2 in. by 8 in. by 16 in. B. 3 in. by 8 in. by 16 in. C. 4 in. by 8 in. by 12 in. D. 3 in. by 8 in. by 14 in.

OpenStudy (bloomlocke367):

well that's pretty easy since it's multiple choice. just multiply each of them out to see which one works (and no, I didn't get the testimonial)

OpenStudy (anonymous):

oh

OpenStudy (bloomlocke367):

did you find the answer?

OpenStudy (bloomlocke367):

2*8*16=? 3*8*16=? 4*8*12=? 3*8*14=?

OpenStudy (anonymous):

oh D silly me :P

OpenStudy (anonymous):

100% :D your the best

OpenStudy (anonymous):

i thought still have a few more hold on

OpenStudy (bloomlocke367):

thanks :) I'm here to help

OpenStudy (bloomlocke367):

open a new question :)